The coil of an electric bulb takes 40 watts to start glowing. If more than 40 W is supplied, 60% of the extra power is converted into light and the remaining into heat. The bulb consumes 100 W at 220 V. Find the percentage drop in the light intensity at a if the supply voltage changes from 220 V to 200 V.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
xSol. P = 100 w V = 220 v Case l : Excess power = 100 – 40 = 60 ww Power converted to light = 60 * 60/100 = 36 w Case ll : Power = (220)^2/484 = 82.64 w Excess power = 82.64 – 40 = 42.64 w Power converted to light = 42.64 * 60/100 = 25.584 w ∆P = 36 – 25.584 = 10.416 Required % = 10.416/36 *100 = 28.93 = 29%